# Automatic Hardware Pragma Insertion in High-Level Synthesis: A Non-Linear Programming Approach

### v1

@@ 7.5. Scalability
To mitigate prolonged solving times for specific kernels and problem sizes, we’ve implemented a 30-minute timeout constraint for the BARON solver. While this timeout doesn’t guarantee achieving optimality, it ensures that the solver provides the best solution it has found within the time limit. In Table 3, we present statistics regarding the number of problem timeouts (ND T/O) and problem non-timeouts (ND NT/O), along with the average time in seconds (Avg Time) for all problems and exclusively for those that did not time out. We can note that the 20 NLP problems for CNN finish in few seconds with an average of 3.71 seconds.
@@
Table 3. Study of the scalability
@@ 8. Related Work
NLP-DSE makes it possible to automatically introduce pragmas in order to obtain a design with a good QoR. Many previous works using different DSE methods have the same objective. Model-free DSEs (Sohrabizadeh et al., 2021; Yu et al., 2018; Ferretti et al., 2018) evaluate each candidate by generating the HLS report. The synthesis time or report generation can last several hours, which considerably reduces the explored design space. Additionally, some DSE, such as (Sohrabizadeh et al., 2021), may lead to local minima. In order to avoid the synthesis time limitation, model-based DSEs and AI-driven DSEs have been developed. They estimate the performance of each design using a cost model (Zhao et al., 2017; Bannwart Perina et al., 2019; Zhong et al., 2016; Wang et al., 2017, 2017; Zhong et al., 2017; Zuo et al., 2015) , Graph Neural Networks (GNN) (Sohrabizadeh et al., 2022; Wu et al., 2023) or decision trees (O’Neal et al., 2018; Yu et al., 2021; Makrani et al., 2019). These DSEs estimate the QoR of the design in a few milliseconds. However, the evaluation of a large number of designs can still take several hours and the accuracy may not match the HLS report. Hence, evaluating only the top-$n$ results with HLS may lead to sub-optimal solutions. Other works allow one-shot optimization with code transformations, e.g., (Ye et al., 2022) but the space of hardware directives and code transformations is limited. Some focus on a predefined micro-architectures, e.g., (Wang et al., 2021) but are limited in their application. NLP-DSE is a hybrid approach: our cost model in the form of NLP makes it possible to consider very large design spaces in a few minutes. This makes it potentially even faster than existing cost models. Yet we rely on HLS to ensure accurate performance.
Recent advances in optimization solvers such as BARON (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) have allowed NLP-based approaches to become a promising alternative to approximate ILP-based methods, as they can encode more complex and realistic performance models. Previous works (Bondhugula et al., 2008; Pouchet et al., 2013a; Zuo et al., 2017) encode a cost model as Linear Programming (LP) problem. However, it requires an approximation, such as approximating the volume of communication inter loops (Bondhugula et al., 2008; Pouchet et al., 2013a), or simplifying the space by exposing direct parallelization in the problem (Zuo et al., 2017).
@@
The selection of tile sizes remains fundamental for the final QoR Similar to our approach, (Liu et al., 2017) uses a cost model to select the tile size. Although their space is much more complete than ours as we are restricted to Merlin’s transformations, their method does not allow the evaluation of the whole space.

### v2

@@ 7.6. Scalability
To mitigate prolonged solving times for specific kernels and problem sizes, we have implemented a 30-minute timeout constraint for the AMPL BARON solver. While this timeout does not guarantee achieving optimality, it ensures that the solver provides the best solution it has found within the time limit. In Table 7, we present statistics regarding the number of problem timeouts (ND T/O) and problem non-timeouts (ND NT/O), along with the average time in seconds (Avg Time) for all problems and exclusively for those that did not time out. We can note that the 20 NLP problems for CNN finish in few seconds with an average of 3.71 seconds.
@@
Table 7. Study of the scalability of the NLP solver across various sizes of Polybench and CNN. Comparison of the number of designs that timeout (ND T/O), the number of designs that do not timeout (ND NT/O), the average time to solve the problem (Avg Time), and the average time to solve the problem for non-timeout designs (Avg Time NT/O).
8. Examples
In this section, we illustrate the significance of our method by contrasting it with AutoDSE and highlighting the advantages of our Design Space Exploration (DSE) approach. Our method excels in addressing domain-specific constraints, providing superior convergence in complex scenarios compared to AutoDSE. Furthermore, our DSE demonstrates adaptability and efficiency, proving to be robust in handling intricate design spaces, offering a more versatile and high-performing solution.
8.1. 2mm Medium
1#pragma ACCEL PIPELINE PIPE_L0
2#pragma ACCEL TILE FACTOR=TILE_L0
3#pragma ACCEL PARALLEL FACTOR=PARA_L0
4 Loop0: for (i1 = 0; i1 < 180; i1++) {
5#pragma ACCEL PIPELINE PIPE_L2
6#pragma ACCEL TILE FACTOR=TILE_L2
7#pragma ACCEL PARALLEL FACTOR=PARA_L2
8 Loop1: for (j1 = 0; j1 < 190; j1++) {
9 S0: tmp[i1][j1] = 0.0;
10#pragma ACCEL PARALLEL FACTOR=PARA_L4
11 Loop2: for (k1 = 0; k1 < 210; ++k1) {
12 S1: tmp[i1][j1] += alpha * A[i1][k1] * B[k1][j1];
13 }
14 }
15 }
16#pragma ACCEL PIPELINE PIPE_L1
17#pragma ACCEL TILE FACTOR=TILE_L1
18#pragma ACCEL PARALLEL FACTOR=PARA_L1
19 Loop3: for (i2 = 0; i2 < 180; i2++) {
20#pragma ACCEL PIPELINE PIPE_L3
21#pragma ACCEL TILE FACTOR=TILE_L3
22#pragma ACCEL PARALLEL FACTOR=PARA_L3
23 Loop4: for (j2 = 0; j2 < 220; j2++) {
24 S2: D[i2][j2] *= beta;
25#pragma ACCEL PARALLEL FACTOR=PARA_L5
26 Loop5: for (k2 = 0; k2 < 190; ++k2) {
27 S3: D[i2][j2] += tmp[i2][k2] * C[k2][j2];
28 }
29 }
30 }
31 }
32}
Listing 11: Implementing the 2mm code with pragma directives for pipelining, tiling, and parallelization for each loop: $D=\alpha\times A\times B\times C+\beta\times D$.
2mm serves as a linear algebra kernel, acting as a surrogate for transformer inference, such as Bert. The code snippet in Listing 11 illustrates the Medium-sized configuration with potential pragma options. The PIPE pragma can be either flattened or off (default), while PARA and TILE can be any divisor of the loop trip count, defaulting to 1.
8.1.1. AutoDSE
The optimal design identified by AutoDSE involves: PARA_L5 = 220, PIPE_L3 = flatten, PARA_L4 = 4 with all other parameters set to 1 or off. However, AutoDSE faces challenges in achieving a high Quality of Results (QoR) for 2mm due to two primary reasons:
For 2mm AutoDSE does not allow you to find a design with good QoR for two main reasons:
  • three out of four of workers initially over-utilize parallelism by flattening PIPE_L0 and $\setminus$ or PIPE_L1 (and hence unroll the innermost loops), causing timeouts in current High-Level Synthesis (HLS) tools. Even without considering timeouts, these designs exceed array partitioning limits, preventing the reading of all data in a single cycle as expected by the unrolling process. Moreover, these designs strain hardware resources, requiring backtracking, which extends the search duration.
  • one out of four mainly optimize a single loop body and is not able to optimize the second loop body. Moreover, even the loop body optimizer is not perfectly optimized and can be more parallelized.
8.1.2. NLP-DSE
Now, we delve into how NLP-DSE overcomes these challenges to discover designs with superior QoR, emphasizing the usefulness of the Design Space Exploration (DSE) described in section 6.
The initial NLP-DSE design features parameters: PARA_L0 = 3, PIPE_L2 = flatten, PARA_L4 = 210, PARA_L1 = 6, PIPE_L3 = flatten, PARA_L5 = 190 achieving 13 GFLOPS/s. However, the compiler fails to apply PARA_L0 and PARA_L1 pragmas, creating a performance gap.
The second design is with the parameters: PIPE_L2 = flatten, PARA_L2 = 2, PARA_L4 = 210, PARA_L3 = 5, PIPE_L3 = flatten, PARA_L5 = 190 which allows to achieve a throughput of 85 GFLOPS/s. However the unroll factor of PARA_L3 does not allow the compiler Xilinx Merlin to transfer the data from off-chip to on-chip efficiently which does not allow to achieve the lower bound. Hence we continue the search and we found our the design with the best QoR in our space at the iteration 8 with the parameters: PIPE_L2 = flatten, PARA_L2 = 2, PARA_L4 = 210, PARA_L3 = 2, PIPE_L3 = flatten, PARA_L5 = 190 We can note that Xilinx Merlin did not allow the data to be transferred optimally so we did not achieve the lower bound and we continue the search. From iteration 10 the lower bound found by the NLP is greater than the latency (HLS report) of design 8, which makes it possible to stop the search because even if we reach the lower bound we will have a latency greater than what we have already obtained.
The Figure 6 summarize the result achieved at each step of the DSE. The step 4,5 and 6 (red) found the same configurations as the step 2 and 3 so we do not need to run the synthesis as we already have the results.
We execute our Design Space Exploration (DSE) using 8 threads, allowing us to concurrently evaluate multiple designs in parallel. This approach anticipates that certain designs may not achieve the desired performance, and by running 8 designs simultaneously, we efficiently explore the design space. For this specific example, we perform a single iteration of the DSE.
Figure 6. Representation of the throughput (GF/s) achieved for each design obtained at each stage of the NLP-DSE for the 2mm kernel.
@@ 9. Related Work
NLP-DSE makes it possible to automatically introduce pragmas in order to obtain a design with a good QoR. Many previous works using different DSE methods have the same objective. These model-free Design Space Exploration (DSE) techniques, as exemplified by works such as (Sohrabizadeh et al., 2021; Yu et al., 2018; Ferretti et al., 2018), employ a methodology where the compiler acts as a black-box, and they dynamically adapt their exploration strategies based on the outcomes of previous iterations. In these approaches, each candidate design is evaluated by generating a High-Level Synthesis (HLS) report. However, the time required for the synthesis or report generation can extend over several hours, significantly limiting the breadth of the explored design space. Moreover, it is worth noting that certain DSE methodologies, including those described in (Sohrabizadeh et al., 2021), might encounter challenges such as converging to local minima, which can impede the discovery of the globally optimal solution.
To circumvent the constraints imposed by synthesis time, novel approaches in Design Space Exploration (DSE) have emerged, including model-based DSEs and AI-driven DSEs. These methods leverage sophisticated techniques such as cost modeling (Zhao et al., 2017; Bannwart Perina et al., 2019; Zhong et al., 2016; Zuo et al., 2015a), Neural Networks (NN) (Wang et al., 2017; Zhong et al., 2017; Koeplinger et al., 2016; Liu and Carloni, 2013; Liu et al., 2019; Schafer and Wakabayashi, 2012; Xydis et al., 2015), Graph Neural Networks (GNN) (Sohrabizadeh et al., 2023b, 2022; Wu et al., 2023), or decision trees (DT) (O’Neal et al., 2018; Yu et al., 2021; Makrani et al., 2019) to estimate the Quality of Results (QoR) of each design rapidly. By utilizing these techniques, the evaluation time for a single design can be reduced to mere milliseconds. However, despite this acceleration, assessing a large number of designs still entails a significant time investment. Furthermore, while these rapid evaluations provide valuable insights, they may not perfectly align with the outcomes obtained from High-Level Synthesis (HLS) reports in terms of accuracy. Consequently, relying solely on HLS validation for the top-$n$ results may lead to suboptimal solutions, as the rapid evaluation methods might not capture all pertinent design intricacies. Therefore, a more comprehensive approach that combines the strengths of both rapid evaluation techniques and traditional HLS validation is necessary to ensure optimal design outcomes.
In contrast, alternative methodologies offer one-shot optimization through code transformations and pragma insertion, as evidenced by works like (Ye et al., 2022; Lai et al., 2019; Huang et al., 2021). However, the efficacy of these approaches is constrained by the limited scope of available hardware directives and code transformations. While some endeavors concentrate on predefined micro-architectures, as seen in (Wang et al., 2021; Basalama et al., 2023b), their applicability is restricted. Moreover, specialized applications such as Deep Neural Networks (DNN) (Zhang et al., 2018, 2020), stencil computations (Chi et al., 2018), sparse linear algebra operations (Du et al., 2022), and neural networks (Basalama et al., 2023a; Sohrabizadeh et al., 2020), including Convolutional Neural Networks (CNNs) (Reggiani et al., 2019), have garnered attention. Yet, these methods encounter challenges when extrapolated beyond their designated domains, rendering generalizations difficult. Hence, while these approaches offer streamlined optimization strategies and tailored solutions for specific problem domains, their broader applicability beyond their respective niches remains a challenge.
NLP-DSE presents a hybrid methodology, leveraging a NLP-based cost model to swiftly explore expansive design spaces within minutes, potentially outpacing existing models in speed. Nevertheless, to ensure precise performance evaluation, reliance on High-Level Synthesis (HLS) remains integral to our approach. Recent advances in optimization solvers such as BARON (Tawarmalani and Sahinidis, 2005; Sahinidis, 2017) have allowed NLP-based approaches to become a promising alternative to approximate ILP-based methods, as they can encode more complex and realistic performance models. Unlike prior works (Bondhugula et al., 2008; Pouchet et al., 2013b; Zuo et al., 2017) that frame the cost model as Linear Programming (LP) problems, NLP-based methods can handle more complex constraints without necessitating approximations, such as estimating communication volumes across loops (Bondhugula et al., 2008; Pouchet et al., 2013b) or simplifying the space by exposing direct parallelization in the problem (Zuo et al., 2017). It is noteworthy that the comparison landscape lacked other NLP-based methodologies, while Linear Programming methods were deemed less suitable due to their incapacity to manage nonlinear constraints effectively. The inclusion of multiple product terms within the objective function and constraints, such as the product of Unroll Factors (UF) for perfectly nested loops, underscores the necessity for accurate modeling, which is challenging to approximate with linear approximations.
@@
The selection of tile sizes remains fundamental for the final QoR. Similar to our approach, (Liu et al., 2017) uses a cost model to select the tile size. Although their space is much more complete than ours as we are restricted to Merlin’s transformations, their method does not allow the evaluation of the whole space.
Approaches that do not rely on precise static analysis, such as (Sohrabizadeh et al., 2021; Yu et al., 2018; Ferretti et al., 2018; Sohrabizadeh et al., 2022; Wu et al., 2023), can take as input any C/C++ kernel supported by the HLS compiler, thus expanding their applicability to a broader spectrum of programs. In contrast, NLP-DSE , akin to other model-based Design Space Exploration (DSE) approaches (Zhao et al., 2017; Bannwart Perina et al., 2019; Zhong et al., 2016; Zuo et al., 2015a), is constrained to affine programs to ensure precise analysis and facilitate the modeling of latency and resource utilization. While this encompasses a significant subset of programs, including AI kernels, and aligns with the MLIR affine dialect, it does not match the versatility of other frameworks. Consequently, there exists a tradeoff between accuracy and generality. Opting for a subset of programs provides more detailed information, thereby accelerating DSE and potentially improving the QoR.